# Worth Their Weight: Randomized and Regularized Block Kaczmarz Algorithms without Preprocessing

### v2

@@ Appendix E Sampling from a Determinantal Point Process
In Section 4, we have presented the convergence of ReBlocK to a weighted least-squares solution under the uniform sampling distribution $\rho=\mathbf{U}(m,k)$. However, the explicit convergence rate $1-\alpha$ is still unknown. Motivated by recent work on Kaczmarz algorithms with determinantal point process (DPP) sampling [15], we now show that under DPP sampling, ReBlocK converges to the ordinary least-squares solution and the convergence of ReBlocK can have a much better dependence on the singular values of $A$ than mSGD.
@@
Our analysis is similar to parts of the nearly concurrent work [16], though [16] does not incorporate a fixed size $k$ for their regularized DPP distribution..
@@
(93) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{*}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{*}\right\rVert^{2}\\ +\frac{1}{2\lambda\alpha^{2}(T-T_{b})}\cdot\max_{i}(r_{i}^{2}).$
Moreover, for any $\ell<k$ the convergence parameter $\alpha$ satisfies
(94) $\alpha\geq\frac{k-\ell}{(k-\ell)+\kappa_{\ell}^{2}(A)+(m+k-2\ell)\frac{\lambda}{\sigma_{n_{r}}^{2}}},$
@@
(95) $\begin{split}\alpha&\geq\frac{k}{k+2\kappa_{k/2}^{2}(A)+2m\lambda/\sigma_{n}^{2}}\\ &\approx\frac{k}{2\kappa^{2}_{k/2}(A)+2\lambda\kappa_{\rm dem}^{2}(A)}.\end{split}$
As noted in Section 3.1, the convergence parameter for mSGD is bounded by
@@
Hence, when the singular values of $A$ decay rapidly and $\lambda$ is significantly smaller than one, ReBlocK with DPP sampling can have a much faster convergence rate than mSGD.
Our bound for the variance of the tail-averaged estimator is crude and could likely be improved using techniques such as those of [16]. Nonetheless, it is interesting to note that the dependence of the variance on the residual vector can be worse in the case of DPP sampling than in the case of uniform samping. This is because the DPP distribution can potentially sample the residual vector in unfavorable ways. It is thus unclear whether DPP sampling is a good strategy for arbitrary inconsistent systems, even if it can be implemented efficiently
@@
The majority of the proof follows similarly to the proofs of Theorems 3.1 and 4.1, and we only highlight the differences, of which there are three: the fact that $x^{(\rho)}=x^{*}$, the value of the parameter $\alpha$, and the variance term in the tail-averaged bound.
We begin by verifying that $x^{(\rho)}=x^{*}$. For a vector $u\in\mathbb{R}^{m}$, we denote its elementary symmetric polynomial by $p_{k}(u):=\sum_{S\in\binom{[m]}{k}}\prod_{i\in S}u_{i}$, where $\binom{[m]}{k}$ is the set of all subsets of size $k$ from the set $[m]=\{1,\ldots,m\}$. Applying equation (5.3) of [15] by using $B=AA^{\top}+\lambda I$ leads to
(96) $\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(I_{S}(AA^{\top}+\lambda I)I_{S}^{\top})^{-1}I_{S}\right]=\frac{U{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))U^{\top}}{p_{k}(q)},$
where $A=U\Sigma V^{\top}$ denotes the compact singular value decomposition of $A$, and $q_{i}=\sigma_{i}^{2}+\lambda$ when $i\leq n_{r}$ and $q_{i}=\lambda$ otherwise, with $\sigma_{1}\geq\dots\geq\sigma_{n_{r}}$ representing the sorted singular values and $n_{r}$ the rank of $A$. Denote by $\overline{W}=UDU^{\top}$ with the diagonal matrix $D:=\frac{{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))}{p_{k}(q)}\succ 0$. Then, by the definition $x^{(\rho)}={\rm argmin}_{x\in\mathbb{R}^{n}}\|Ax-b\|_{\overline{W}}^{2}$, we have
$A^{\top}\overline{W}Ax^{(\rho)}=A^{\top}\overline{W}b.$
Note that $A^{\top}\overline{W}A=V\Sigma D\Sigma V^{\top}$ and $A^{\top}\overline{W}=V\Sigma DU^{\top}b$. Thus, we can apply $V^{\top}$ on the left and use $D\succ 0$ to conclude
$\Sigma^{2}V^{\top}x^{(\rho)}=\Sigma U^{\top}b.$
@@
We now follow closely the logic of the proof of Lemma 4.1 of [15], making appropriate modifications to handle the fact that in GRK we invert $AA^{\top}+\lambda I$ rather than just $AA^{\top}$. Note also that through (97) we have explicitly verified that $\operatorname{range}(\overline{P})=\operatorname{range}(A^{T})$, which is the needed condition for the contraction properties to hold in our convergence bound.
@@
(98) $B_{\ell}=U{\rm diag}\left(q_{1},\ldots q_{m}\right)U^{\top}+\frac{1}{k-\ell}\sum_{j>\ell}^{m}q_{j}I.$
Note that we have used a mildly simpler and looser approximation than [15] by replacing $\frac{k-\ell-1}{k-\ell}$ with $1$ in the first term.
The same arguments as in [15] then imply that
@@
$=\frac{\sigma_{n_{r}}^{2}}{\sigma_{n_{r}}^{2}+\frac{1}{k-\ell}\sum_{j>\ell}^{n_{r}}\sigma_{j}^{2}+\frac{m+k-2\ell}{k-\ell}\lambda}$
$=\frac{k-\ell}{(k-\ell)+\kappa_{\ell}^{2}(A)+(m+k-2\ell)\frac{\lambda}{\sigma_{n_{r}}^{2}}}.$
@@
Finally, we consider the variance term in the tail-averaged bound. Following the proofs of Lemmas C.1 and 4.1, we can obtain a variance term of
(100) $\frac{1}{2\lambda k\alpha^{2}(T-T_{b})}\mathbb{E}\,_{S\sim\rho}\left\lVert r_{S}\right\rVert^{2}$
which is esentially the same as for ReBlocK-U but with the ordinary least-squares residual $r$ instead of the weighted residual $r^{(\rho)}$. Unfortunately, in the case of DPP sampling, we have no guarantee on how $S$ samples the residual vector $r^{(\rho)}$. Thus, the best we can do is uniformly bound $\left\lVert r_{S}\right\rVert^{2}\leq k\cdot\max_{i}(r_{i}^{2})$. This yields the bound in the theorem.
Appendix F Details of Numerical Experiments
In this appendix we provide more details on the numerical examples throughout the paper. The code to run the experiments can be found at https://github.com/ggoldsh/block-kaczmarz-without-preprocessing.
F.1 Experiments with random matrices
The experiments in Sections 3.3 and 4.2 differ only in how the matrix $A$ is generated. Thus, we first describe the matrix construction for each case separately, then describe the rest of the set-up in a unified manner.
@@ F.1 Experiments with random matrices
Regarding Section 4.2, the examples in Fig. 3 are constructed as discretizations of underlying continuous problems, which is both a realistic scenario and serves to ensure that $A$ contains many nearly singular blocks. In both cases, we first generate an $n\times n$ matrix $C$. For the left panel we set $C=I$, whereas for the right panel we set $C$ with singular values $\sigma_{i}=1/i^{2}$ and random orthonormal singular vectors. Once $C$ is constructed, we define a set of $n$ functions $f_{1},\ldots,f_{n}$ by
@@
Next, we construct a vector $v$ of $m$ coordinates uniformly spaced across the interval $[-1,1]$, namely
@@
so that column $j$ of $A$ is a dicretized representation of the function $f_{j}$. The condition number of the resulting matrices $A$ are $\kappa(A)\approx 11$ in the left panel and $\kappa(A)\approx 3.6\cdot 10^{4}$ in the right panel.
For all experiments in Sections 3.3 and 4.2 we set $m=10^{5}$, $n=10^{2}$, and $k=30$. Furthermore, the vectors $b$ are constructed by setting
@@
where $y\sim\mathcal{N}(0,I_{n})$ and $z_{i}\sim\mathcal{N}(0,10^{-4})$. The initial guess is $x_{0}=0$ and the number of iterations is $T=10^{4}$, which is equivalent to three passes over the data. The step size for minibatch SGD is constant within each run and has been tuned independently for each example, specifically to be as large as possible without introducing any signs of instability, up to a factor of $2$. The value of $\lambda$ for ReBlocK has been set to $\lambda=0.001$ and is not optimized on a per-example basis. The reported quantity is the suboptimality of the solution for the unweighted problem (1), namely $\epsilon=\left\lVert Ax-b\right\rVert/\left\lVert Ax^{*}-b\right\rVert-1=\left\lVert Ax-b\right\rVert/\left\lVert r\right\rVert-1$. Since it is expensive to calculate the residual, this quantity is only calculated and reported every $10$ iterations. These experiments with random matrices are carried out in double precision on a 1.1 GHz Quad-Core Intel Core i5 CPU.
F.2 Natural Gradient Experiments
@@
The neural network model is a simple periodic ResNet [21] with 5 layers. For input $s\in\mathbb{R}$, the network outputs $f(s)\in\mathbb{R}$ given by
@@
The intermediate layers have dimensions $y_{i}\in\mathbb{R}^{500}$ and the weight matrices $W_{i}$ and bias vectors $b_{i}$ have the appropriate dimensions to match. The weights are initialized using a Lecun normal initialization and the biases are all initialized to zero. The parameters are collected into a single vector $\theta\in\mathbb{R}^{753,001}$ for convenience, leading to the neural network function $f_{\theta}(r)$.
The loss function is defined as in (27) and the network is trained using subsampled natural gradient descent, as described for example in [38], with a batch size of $N_{b}=500$, a Tikhonov regularization of $\lambda=0.01$, and a step size of $\eta=0.5$. This corresponds to the parameter update
@@
From this training run, three snapshots are taken and used to generate the least-squares problems for Fig. 5, following Eq. 32. The first snapshot is from the “pre-descent” phase before the loss begins to decrease, the second is from the “descent” phase during which the loss decreases rapidly, and the third is from the “post-descent” phase when the decay rate of the loss has slowed substantially. The snapshots are indicated by the vertical dotted lines in Fig. 10 and form the basis of the experiments of Figs. 5, 6, 7, and 8. The batch size of $k=50$ used in Fig. 5 is meant to represent the realistic scenario when each iteration uses too few samples to thoroughly represent the target function. Furthermore, the continuous problems are treated directly by uniformly sampling $k$ points from the domain $[0,1]$ at each iteration and calculating the network outputs and gradients at these points.

### v3

@@ Appendix E Sampling from a Determinantal Point Process
In Section 5, we have presented the convergence of ReBlocK to a weighted least-squares solution under the uniform sampling distribution $\rho=\mathbf{U}(m,k)$. However, the explicit convergence rate $1-\alpha$ is still unknown. Motivated by recent work on Kaczmarz algorithms with determinantal point process (DPP) sampling [derezinski2024solving], we now show that under DPP sampling, ReBlocK converges to the ordinary least-squares solution and the convergence of ReBlocK can have a much better dependence on the singular values of $A$ than mSGD.
@@
Our analysis is similar to parts of the nearly concurrent work [derezinski2025], though [derezinski2025] does not incorporate a fixed size $k$ for their regularized DPP distribution.
@@
(91) $\mathbb{E}\,\left\lVert\overline{x}_{T}-x^{*}\right\rVert^{2}\leq 2(1-\alpha)^{T_{b}+1}\left\lVert x_{0}-x^{*}\right\rVert^{2}\\ +\frac{4}{\alpha^{2}(T-T_{b})}V$
with $V=\mathbb{E}\,_{S\sim\rho}\left\lVert A_{S}^{\top}(A_{S}A_{S}^{\top}+\lambda kI)^{-1}r_{S}\right\rVert^{2}$. The variance term is bounded by
(92) $V\leq\frac{1}{4\lambda}\cdot\max_{i}r_{i}^{2}.$
Finally, for any $\ell<k$ the convergence parameter $\alpha$ satisfies
(93) $\alpha\geq\frac{k-\ell}{(k-\ell)+\kappa_{\ell}^{2}(A)+(m+k-2\ell)\frac{\lambda k}{\sigma_{n_{r}}^{2}}},$
@@
(94) $\begin{split}\alpha&\geq\frac{k}{k+2\kappa_{k/2}^{2}(A)+2m\lambda k/\sigma_{n}^{2}}\\ &\approx\frac{k}{2\kappa^{2}_{k/2}(A)+2\lambda k\kappa_{\rm dem}^{2}(A)}.\end{split}$
As noted in Section 4, the convergence parameter for mSGD is bounded by
@@
Hence, when the singular values of $A$ decay rapidly and $\lambda k$ is significantly smaller than one, ReBlocK with DPP sampling can have a much faster convergence rate than mSGD.
Our bound for the variance of the tail-averaged estimator is crude and could likely be improved using techniques such as those of [derezinski2025]. Nonetheless, it is interesting to note that the dependence of the variance on the residual vector can be worse in the case of DPP sampling than in the case of uniform sampling. This is because the DPP distribution can potentially sample the residual vector in unfavorable ways. It is thus unclear whether DPP sampling is a good strategy for arbitrary inconsistent systems, even if it can be implemented efficiently
@@
The majority of the proof follows similarly to the proofs of Theorems 3.1 and 5.1, and we only highlight the differences, of which there are three: the fact that $x^{(\rho)}=x^{*}$, the value of the parameter $\alpha$, and the bound on the variance term in the tail-averaged bound.
We begin by verifying that $x^{(\rho)}=x^{*}$. For a vector $u\in\mathbb{R}^{m}$, we denote its elementary symmetric polynomial by $p_{k}(u):=\sum_{S\in\binom{[m]}{k}}\prod_{i\in S}u_{i}$, where $\binom{[m]}{k}$ is the set of all subsets of size $k$ from the set $m]=\{1,\ldots,m\}$. Applying equation (5.3) of [derezinski2024solving] by using $B=AA^{\top}+\lambda kI$ leads to
(95) $\overline{W}=\mathbb{E}_{S\sim\rho}\left[I_{S}^{\top}(I_{S}(AA^{\top}+\lambda kI)I_{S}^{\top})^{-1}I_{S}\right]=\frac{U{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))U^{\top}}{p_{k}(q)},$
where $A=U\Sigma V^{\top}$ denotes the full singular value decomposition of $A$, and $q_{i}=\sigma_{i}^{2}+\lambda k$ when $i\leq n_{r}$ and $q_{i}=\lambda k$ otherwise, with $\sigma_{1}\geq\dots\geq\sigma_{n_{r}}$ representing the sorted singular values and $n_{r}$ the rank of $A$. Denote by $\overline{W}=UDU^{\top}$ with the diagonal matrix $D:=\frac{{\rm diag}(p_{k-1}(q_{-1}),\ldots,p_{k-1}(q_{-m}))}{p_{k}(q)}\succ 0$. Then, by the definition $x^{(\rho)}={\rm argmin}_{x\in\mathbb{R}^{n}}\|Ax-b\|_{\overline{W}}^{2}$, we have
(96) $A^{\top}\overline{W}Ax^{(\rho)}=A^{\top}\overline{W}b.$
Note that $A^{\top}\overline{W}A=V\Sigma^{\top}D\Sigma V^{\top}$ and $A^{\top}\overline{W}=V\Sigma^{\top}DU^{\top}b$. Also, because $\Sigma\in\mathbb{R}^{m\times n}$ and $D\in\mathbb{R}^{m\times m}$ are diagonal, we have $\Sigma^{\top}D=\hat{D}\Sigma^{\top}$, where $\hat{D}:=D[1:n,1:n]$ is the leading principal submatrix of $D$. Substituting this into (96) gives $V\hat{D}\Sigma^{\top}\Sigma V^{\top}x^{(\rho)}=V\hat{D}\Sigma^{\top}U^{\top}b$. Thus, we can apply $V^{\top}$ on the left and use $D\succ 0$ to conclude
$\Sigma^{\top}\Sigma V^{\top}x^{(\rho)}=\Sigma^{\top}U^{\top}b.$
@@
We now follow closely the logic of the proof of Lemma 4.1 of [derezinski2024solving], making appropriate modifications to handle the fact that in ReBlocK we invert $AA^{\top}+\lambda kI$ rather than just $AA^{\top}$. Note also that through (97) we have explicitly verified that $\operatorname{range}(\overline{P})=\operatorname{range}(A^{T})$, which is the needed condition for the contraction properties to hold in our convergence bound.
@@
(98) $B_{\ell}=U{\rm diag}\left(q_{1},\ldots q_{m}\right)U^{\top}+\frac{1}{k-\ell}\left(\sum_{j>\ell}^{m}q_{j}\right)I.$
Note that we have used a mildly simpler and looser approximation than [derezinski2024solving] by replacing $\frac{k-\ell-1}{k-\ell}$ with $1$ in the first term.
The same arguments as in [derezinski2024solving] then imply that
@@
$=\frac{\sigma_{n_{r}}^{2}}{\sigma_{n_{r}}^{2}+\frac{1}{k-\ell}\sum_{j>\ell}^{n_{r}}\sigma_{j}^{2}+\frac{m+k-2\ell}{k-\ell}\lambda k}$
$=\frac{k-\ell}{(k-\ell)+\kappa_{\ell}^{2}(A)+(m+k-2\ell)\frac{\lambda k}{\sigma_{n_{r}}^{2}}}.$
@@
Finally, we consider the bound for the variance term in the tail-averaged bound. Following the proof of Theorem 5.1 we can obtain an initial bound of
(100) $V\leq\frac{1}{4\lambda k}\cdot\mathbb{E}\,_{S\sim\rho}\left\lVert r_{S}\right\rVert^{2}.$
Unfortunately, in the case of DPP sampling, we have no guarantee on how $S$ samples the residual vector $r^{(\rho)}$. Thus, the best we can do is uniformly bound $\left\lVert r_{S}\right\rVert^{2}\leq k\cdot\max_{i}r_{i}^{2}$. This yields the bound in the theorem.
In this supplement we provide more details on the numerical examples throughout the paper. The code to run the experiments can be found at https://github.com/ggoldsh/block-kaczmarz-without-preprocessing.
Appendix F Experiments with random matrices
In this section we provide the details of the synthetic experiments from Sections 3.1, 4.1, and 5.1. All of these experiments have $m=10^{5}$, $n=10^{2}$, and $k=30$. The experiments differ only in how the matrix $A$ is generated. Thus, we first describe the matrix construction for each case separately, then describe the rest of the set-up in a unified manner.
@@ Appendix F Experiments with random matrices
Regarding Sections 3.1 and 5.1, the examples these sections are constructed as discretizations of underlying continuous problems, which is both a realistic scenario and serves to ensure that $A$ contains many nearly singular blocks. In both cases, we first generate an $n\times n$ matrix $C$. For the left panel we set $C=I$, whereas for the right panel we set $C$ with singular values $\sigma_{i}=1/i$ and random orthonormal singular vectors. Once $C$ is constructed, we define a set of $n$ functions $f_{1},\ldots,f_{n}$ by
@@
Next, we construct a vector $v$ of $m$ coordinates uniformly spaced across the interval [1,1][-1,1], namely
@@
so that column $j$ of $A$ is a discretized representation of the function $f_{j}$. The condition number of the resulting matrices $A$ are $\kappa(A)\approx 11$ in the left panel and $\kappa(A)\approx 450$ in the right panel.
For all experiments in Sections 3.1, 4.1, and 5.1 the vectors $b$ are constructed by setting
@@
where $y\sim\mathcal{N}(0,I_{n})$ and $z_{i}\sim\mathcal{N}(0,\sigma)$. The noise level is set to $\sigma=1e-2$ except for in Figure 7 where the noise level is explicitly varied. The initial guess is $x_{0}=0$ and the number of iterations is $T=10^{5}$, which is equivalent to thirty passes over the data. The step size for minibatch SGD is constant within each run and has been tuned independently for each example, specifically to be as large as possible without introducing any signs of instability, up to a factor of $2$. The value of $\lambda$ for ReBlocK has been set to $\lambda=0.001$ and is not optimized on a per-example basis. The reported quantity is the relative solution error $\left\lVert x-x^{*}\right\rVert/\left\lVert x^{*}\right\rVert$. These experiments with random matrices are carried out in double precision on a 1.1 GHz Quad-Core Intel Core i5 CPU.
Appendix G Natural Gradient Experiments
@@
The neural network model is a simple periodic ResNet [he2016deep] with 5 layers. For input $s\in\mathbb{R}$, the network outputs $f(s)\in\mathbb{R}$ are given by
@@
The intermediate layers have dimensions $y_{i}\in\mathbb{R}^{500}$ and the weight matrices $W_{i}$ and bias vectors $b_{i}$ have the appropriate dimensions to match. The weights are initialized using a Lecun normal initialization and the biases are all initialized to zero. The parameters are collected into a single vector $\theta\in\mathbb{R}^{753,001}$ for convenience, leading to the neural network function $f_{\theta}(s)$.
The loss function is defined as in (29) and the network is trained using subsampled natural gradient descent, as described for example in [ren2019efficient], with a batch size of $N_{b}=500$, a Tikhonov regularization of $\lambda=0.01$, and a step size of $\eta=0.5$. This corresponds to the parameter update
@@
From this training run, three snapshots are taken and used to generate the least-squares problems for Figs. 8, 9, 10, 11, and 12, following Eq. 34. The first snapshot is from the “pre-descent” phase before the loss begins to decrease, the second is from the “descent” phase during which the loss decreases rapidly, and the third is from the “post-descent” phase when the decay rate of the loss has slowed substantially. The snapshots are indicated by the vertical dotted lines in Fig. 14 and form the basis of the experiments of Figs. 8, 9, 10, 11, and 12. The batch size of $k=50$ used in Fig. 8 is meant to represent the realistic scenario when each iteration uses too few samples to thoroughly represent the target function. Furthermore, the continuous problems are treated directly by uniformly sampling $k$ points from the domain $[0,1]$ at each iteration and calculating the network outputs and gradients at these points.